OpenStudy (jkbo):

please help Solve the equation -12x − 8 = -7x − 4

OpenStudy (compassionate):

Hi, My name is Stephen and I will be your tutor. -12x − 8 = -7x − 4 Okay, so the first step is gathering your x's and numerical values to one side. X's go to the left, and numbers to the right. We do this by performing the opposite action. Since -7x is negative, we add it to both sides, and since -8 is negative, we add it to both sides. -12x − 8 = -7x − 4 -12x + 7x = 8 - 4 Okay, so we simply add these two together like so... -5x = 4 Now we divide... x = -4/5 or -0.8

OpenStudy (jkbo):

thank you very much

OpenStudy (jkbo):

@Compassionate can u help with one more ? x − 9 = x − 10 Solution: x = -10

OpenStudy (compassionate):

Last one before I go to bed. x - 9 = x - 10 Add 9. x = x - 1 Subtract x x - x = - 1 0 = - 1 This is a false statement.

OpenStudy (jkbo):

thank you